OpenStudy (jaylelile):

I'm having a hard time remembering how to go about this problem. help? Problem below

OpenStudy (jaylelile):

\[\frac{ 4}{ 8 }^{2}b+\frac{ 40 }{ 32 }^{3}b\]

OpenStudy (jaylelile):

@Nnesha ?

Nnesha (nnesha):

well the variables are the same so they are like terms combine them

Nnesha (nnesha):

here is an example when the variables and the exponents are the same we can add/subtract their coefficients \[\rm 2x^2+5x^2 = (2+5)x^2=7x^2\]

OpenStudy (jaylelile):

\[\frac{ 44}{ 40 }^{3}b ?????\]

Nnesha (nnesha):

hmm we didn't add them correctly :=) \[\rm \frac{ 4}{ 8 }^{2}b+\frac{ 40 }{ 32 }^{3}b=(\frac{4^2}{8}+\frac{40^3}{32})b\]

OpenStudy (jaylelile):

that seems too simple though.. lol

Nnesha (nnesha):

\[\rm \frac{ 4}{ 8 }^{2}b+\frac{ 40 }{ 32 }^{3}b=(\frac{4^2}{8}+\frac{40^3}{32})b\] \[\rm (\color{ReD}{\frac{4^2}{8}+\frac{40^3}{32}})b\] solve paretheses what's the common denominator ??

OpenStudy (jaylelile):

do we add the exponents?

OpenStudy (jaylelile):

I'm so confused

Nnesha (nnesha):

no when we `multiply` same bases then we should add their exponents when we combine like terms base stay the same we just have to add/subtract their `coefficients `

OpenStudy (jaylelile):

I'm still getting \[\frac{ 44 }{ 40 }^{3}b\] ..........

Nnesha (nnesha):

alright show some work i mean what was your step how did you get 40 at the denominaotr ?

Nnesha (nnesha):

i see what you did there common mistake

OpenStudy (jaylelile):

I just added 8 and 32... is that wrong?

Nnesha (nnesha):

\[\frac{ 3 }{ 2} + \frac{4}{6}\] wrong that's not how we should add